OpenStudy (anonymous):

a cylinder and a rectangular prism have the same volume and same height. the base of the prism is a square with side length 8 cm. what is the approxmate radius of the cylinder?

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

let r be the radius and h be the height of the cylinder \[volume~ of~ cylinder=\pi r^2 h\] volume of prism=base area * height =8*8*h=64 h equate the two volumes and find r

OpenStudy (anonymous):

Is it a multiple choice question?